OpenStudy (anonymous):

Somebody!! please help, will medal Solve for t h= 7cos(pi/3t)

OpenStudy (anonymous):

@Hero @jim_thompson5910 @satellite73

OpenStudy (anonymous):

\[t=\pi/(3*\cos^{-1}( (h/7)))\]

OpenStudy (anonymous):

Divide both sides by 7 h/7=cos(pi/3t) cos^-1(h/7)=pi/3t, take the inverse cosine of both sides. 1/(cos^-1(h/7)=3t/pi, find the reciprocal of both sides. pi/(3*cos^-1(h/7)=t, multiple by pi and divide by 3.

OpenStudy (tkhunny):

You will have to clarify the notation. pi/3t means what? Choices might be: \(\dfrac{\pi}{3}t\;or\;\dfrac{\pi}{3t}\) You MUST write clearly and unambiguously.

OpenStudy (anonymous):

Well, that is of fault in the original question as well.

OpenStudy (anonymous):

sorry, its pi/3 t. they are seperate

OpenStudy (tkhunny):

Use parentheses to clarify. Spaces only sometimes help. MAKE SURE it says what you want. Also, keep in mind that not everyone knows how to be clear. Don't just copy what someone else wrote. Textbooks, computers, and humans all have slightly different rules in order to be clear.